the sum of four consecutive whole number is 150
a. let x represent the first number. write expression to represent the other three numbers.
b. write and solve an equation to find the four numbers

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

the sum of four consecutive whole number is 150
a. let x represent the first number. write expression to represent the other three numbers.
1st: x
2nd: x+1
3rd: x+2
4th: x+3
------------
b. write and solve an equation to find the four numbers
-----------------------------------
Add those expressions to get:
4x+6 = 150
4x = 144
x = 38
x+1 = 39
x+2 = 40
x+3 = 41
-----------------------
Cheers,
Stan H.

Answer by Alan3354(69443)

the sum of four consecutive whole number is 150
a. let x represent the first number. write expression to represent the other three numbers.
x+1, x+2, x+3
--------------------------
b. write and solve an equation to find the four numbers
x + x+1 + x+2 + x+3 = 150
4x + 6 = 150
x = 36
--> 36, 37, 38, 39
-------------------
-------------------
If you have time to waste do it as above.
If not:
150/4 = 37.5, the center of the 4 numbers
--> 36, 37, 38, 39